A number when successively divided by 9, 11 and 13 leaves remainders 8, 9 and 8 respectively.Finf the number.


Grade:10

2 Answers

Ashwin Muralidharan IIT Madras
12 years ago

Hi Sathyaram,

 

First when the number is divided by 9. the remainder is 8.

So N = 9x+8.

Similarly, next x = 11y+9, and y=13z+8.

So N = 99y+89 = 99(13z+8)+89 = 1287z+792+89 = 1287z+881.

So N is of the form, 1287*(A whole Number)+881.

 

If you need to find the minimum number, then it would be 881.

 

Hope that helps.

 

All the best.

Regards,

Ashwin (IIT Madras).

Sathya
12 years ago

Thanks for your reply.It was very helpful! Smile